OpenStudy (anonymous):

how would i solve this word problem, i know have to use 2 equations though A kayaker paddled 2 hours with a 6 mph current in a river. The return trip against the same current took 3 hours. Find the speed the kayaker would make in still water.

OpenStudy (anonymous):

Have the kayaker's speed be x. If he's going downstream, his total speed will be x+6. If he's going upstream, his speed will be x-6. To find the distance traveled in terms of x, multiply the speed by the time. d=2x+12 d=3x-18 Since we know that he went the same distance each time, we can set these new equations equal to each other. 2x+12=3x-18 solve for x THIS KINDA CONFUSING.. not sure yet so hold up???

OpenStudy (anonymous):

The key to this problem is that the distance is the same downstream and upstream. From one of our favorite formulas d=rt, we will get\[2(v+6)=3(v-6)\]which is to say that the distance downstream is the same as the distance upstream. Solve for v to get your answer.

OpenStudy (anonymous):

yeahhh @animalain got it!!

OpenStudy (anonymous):

u both got it lol give each other a medal lol

OpenStudy (anonymous):

katie seu bem-vindos!!!! lol!! ;p

OpenStudy (anonymous):

lol what?!?!?! was that an instult lol

OpenStudy (anonymous):

Did you get 30 mph?

OpenStudy (anonymous):

yess i did!

OpenStudy (anonymous):

hahaha insult!! lol!!! of course not!! i said ur welcome kate!! ;p

OpenStudy (anonymous):

kATIE* SRYY

OpenStudy (anonymous):

lol ohhh ok well good ;p and its cool

OpenStudy (anonymous):

wats cool??

OpenStudy (anonymous):

that u spelled my name rong lol its alright

OpenStudy (anonymous):

ohhhhhh true! lol!! ;D

OpenStudy (anonymous):

ppl call me kate all the time :P